Word problem:
A chainsaw runs on fuel that is 10% oil and 90% gas. Howmuch oil will have to be added to 3 gallons of fuel that is 5% oil and 95% gas in order to use the chainsaw?

Answer by ptaylor(2198)

Let x=amount of oil that has to be added
The amount of pure gas before the oil is added (0.95*3) has to equal the amount of pure gas after the oil is added 0.90(3+x), so:
0.95*3=0.90(3+x)
2.85=2.70+0.90x subtract 2.70 from each side
2.85-2.70=2.70-2.70+0.90x collect like terms
0.15=0.90x divide both sides by 0.90
x=0.16667 gal ------------------ans
CK
3*0.95=0.90*(3.16667)
2.85=2.85
Hope this helps---ptaylor